The length of half-wave antenna is at 30 MHz


A. 2.5 meter

B. 5 meter

C. 10 meter

D. 100 meter


1
Expert's answer
2021-01-18T08:12:28-0500

"\begin{aligned}
f &= 30\ MHz = 30 × 10^6Hz \\
&= 3 × 10^7 Hz 
\end{aligned}"

"c = 3 \u00d7 10^8 ms^{-1}"



from,

"c = f\lambda"

"\lambda = \frac{c}f"


"\lambda = \dfrac{3×10^8}{3×10^7} = 10m"


length of half-wave dipole = "\frac{10}{2}m = 5m"

Therefore, the length of half-wave antenna is 5m (B.)